OpenStudy (anonymous):

An absolute value equation of the form |ax+b|=c, where a, b, and c are real numbers, has how many solutions? A) zERO b) oNE c) tWO d) zero, One, or two

OpenStudy (gorv):

if we remove modulus

OpenStudy (gorv):

\[ax+b= \pm c\]

OpenStudy (gorv):

ax+b=c ax+b=-c

OpenStudy (gorv):

ax=c-b x=(c-b)/a and x=(-c-b)/a

OpenStudy (gorv):

two solution it can have